OpenStudy (anonymous):

The equation of the line that goes through the points (−2,6) and (8,5) can be written in general form (with integer coefficients, positive lead coefficient, and no common factor)

OpenStudy (anonymous):

y=-0.1x+5.2

OpenStudy (anonymous):

thats what I got but my system keeps telling me that is the wrong answer

OpenStudy (anonymous):

it's the right answer though...

OpenStudy (phi):

General form http://en.wikipedia.org/wiki/Linear_equation#general%20form

OpenStudy (phi):

Ax + By + C= 0 (with integer coefficients, positive lead coefficient, and no common factor)

OpenStudy (anonymous):

they want it in Ax+By+C=0 form...which i thought would be -0.1x+y+5.2=0

OpenStudy (phi):

but with integer coefficients. multiply both sides by 10

OpenStudy (phi):

isn't it -5.2 ?

OpenStudy (phi):

check that, -5.8

OpenStudy (anonymous):

now im confused...wouldnt the slope be 10

OpenStudy (anonymous):

i meant -1/10...and it says the coefficients have to be positive so thats why im lost

OpenStudy (phi):

you start with 0.1x+y-5.8= 0 (we can check this is correct 0.1*-2+6 -5.8= 0) multiply by 10 x+10y-58=0

OpenStudy (anonymous):

but 0.1 is negative...how did you get positive

OpenStudy (phi):

or from slope-intercept form y= -0.1x +5.8 0.1x + y - 5.8 =0 1x + 10y -58=0 or simply x +10y -58=0 you can check by plugging in (x,y) values that are on the line

OpenStudy (phi):

make sense?

OpenStudy (anonymous):

yes...thank you